• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name meral

    Log entry time 08:02:09 on August 05, 2006

    Entry number 177547

    keyword=Owl Shift/Run Summary

    Owl Shift Summary
    0:00- bmp positions still varying
    0:05- optics target in, 20 uA
    0:10- 4cm dummy target in, 8 uA
    0:20- 15cm dummy target in, 8uA
    0:20- no current
    1:00- current is back
    1:05- Ta target in
    1:30- reboot the ioc HRS-R
    - took the opportunity to change HRS-R p=0.346 GeV/c
    - unsuccessful in changing the dipole current to reflect new momentum
    setting for HRS-R, so this must be fixed
    1:48- mcc having a problem with pss system
    3:00- S0 in, LH2 in, rad out, trigger downloaded, cross section
    4:20- 1uA
    5:30- Requested Hall Access because S0 will not come out!
    6:15- S0 successfully out
    6:30- 4uA requested
    7:11- realized we did not reset the FPP voltage after Hall Access, fixed

    this
    Owl Run Summary
    Kinematics Point 3, p=0.53290 GeV/c, theta=14.0
    Run # Event # Current
    2139 1104460 4uA
    20616 140477 4uA
    2142 758973 8uA
    20620 56526 8uA
    2143 28177 junk
    20621 95427 8uA
    2144 498489 8uA
    2145 47754 8uA
    20622 2600 8uA
    2146 55874 12uA
    20623 30552 12uA
    2147 20210 junk
    20624 8761 junk
    2148 510353 20uA
    20625 215782 20uA
    2149 2001927 2uA
    20626 250191 2uA
    2150 1M 1uA
    20627 167K 1uA
    Kinematics Point 3b p=0.5252 GeV/c, radiator out
    2151 junk 4uA, did not reset the FPP voltage
    20628 junk "
    2152 junk 4uA, did not reset the FPP voltage
    20629 junk "
    2153 2M 4uA
    20630 155K 4uA